A lamp is marked 12V/10W. the lamp is connected to 24V accumulator. in series with lamp you must then connect a resistor.
Calculate the resistance of the resistor!

Answer by Alan3354(69443)

A lamp is marked 12V/10W. the lamp is connected to 24V accumulator. in series with lamp you must then connect a resistor.
Calculate the resistance of the resistor!
-----------------
Since 12V is 1/2 of 24V, it'll be the same as the lamp's resistance.
10W/12V = 5/6 amps
R = 12/(5/6) = 14.4 ohms